# Rhythmic sharing: A bio-inspired paradigm for zero-shot adaptation and learning in neural networks

### v3

@@ Abstract
The brain can rapidly adapt to new contexts and learn from limited data, a coveted characteristic that artificial intelligence algorithms have struggled to mimic. Inspired by oscillatory rhythms of the mechanical structures of neural cells, we developed a learning paradigm that is based on oscillations in link strengths and associates learning with the coordination of these oscillations. We find that this paradigm yields rapid adaptation and learning in artificial neural networks. Link oscillations can rapidly change coordination, endowing the network with the ability to sense subtle context changes in an unsupervised manner. In other words, the network generates the missing contextual tokens required to perform as a generalist AI architecture capable of predicting dynamics in multiple contexts. Oscillations also allow the network to extrapolate dynamics to never-seen-before contexts. These capabilities make our learning paradigm a powerful starting point for novel models of learning and cognition. Furthermore, learning through link coordination is agnostic to the specifics of the neural network architecture, hence our study opens the door for introducing rapid adaptation and learning capabilities into leading AI models.
@@ 1.1 A new learning paradigm
The second part of the paradigm is that the phases of each link can change depending on the information it processes, e.g., different states, as illustrated in fig. 1C. We propose that links that are active in processing a given state can adjust their phases rapidly to synchronize to other links. This synchronization is mediated by objects that can integrate information from subsets of links (inspired by the integrative role of astrocytes and denoted as stars in 1C).
Because input of a different state results in active link changes, state changes manifest as collective changes to synchronization of the link phases as illustrated in 1C. Therefore, link phase synchronization also functions as a classification token of each state.
@@ 2.1.3 Detecting warning signs of catastrophic phenomena
A well-known chaotic system that exhibits this behavior is the 3D Lorenz system[15]. Its internal parameter $\lambda_{\text{Lorenz}}$ is responsible for state changes. The system exhibits a chaotic trajectory when $\lambda_{\text{Lorenz}}=24.5$. When $\lambda_{\text{Lorenz}}$ abruptly changes to $23.5$, the chaotic trajectory shifts into another similar, but transient chaotic trajectory until it suddenly collapses into a dead signal after some probabilistic time[16]. Indeed, fig. 3C shows the close similarity in the spectral properties of the two chaotic trajectories; to resolve the differences in the spectral peaks (0.1-0.2 Hz), at least 50-100 cycles of this frequency band must be observed.
@@ 4.2 Rhythmic sharing
where $\circ$ denotes the Hadamard product. $\bm{\Phi}$ is a vector of phases $[\phi_{1}\,...\,\phi_{N_{l}}]^{\text{T}}$ with $N_{l}$ being the total number of links. $\bm{\omega}_{0}$ is a vector of initially imposed natural frequencies of the links. The vector of local mean fields $\bm{\Psi}(t)\in\mathbb{R}^{N_{l}}$ that each link is coupled to is given by
@@
where $\mathbf{r}\in\mathbb{R}^{N_{l}}$ is a vector of local order parameters whose elements describe the degree of synchronization across the particular local mean field that each link is coupled to. $\mathbf{\hat{A}}_{{\bm{\Phi}}_{ij}}\equiv\mathbf{A}_{{\bm{\Phi}}_{ij}}/|\mathbf{A}_{{\bm{\Phi}}_{i\bullet}}|$, where $\mathbf{A}_{\bm{\Phi}}\in\mathbb{R}^{N_{l}\times N_{l}}$ is a random phase adjacency matrix, binary for simplicity, whose elements equal 1 if the links are connected and 0 otherwise, and the normalization factor $|\mathbf{A}_{{\bm{\Phi}}_{i\bullet}}|$ is the $L^{1}$ norm of the $i$-th row of $\mathbf{A}_{\bm{\Phi}}$. The density of non-zero elements of $\mathbf{A}_{\bm{\Phi}}$ matrix is a tuneable hyperparameter. We note in passing that while we limit the focus of this study to 1:1 synchronization, arbitrary orders of $p:q$ synchronization (for clusters) may be readily implemented by modifying the argument of the right-hand side of the eq. 3 by taking $\sin(\bm{\Psi}-\bm{\Phi})\rightarrow\sin(\mathbf{p}\circ\bm{\Psi}-\mathbf{q}\circ\bm{\Phi})$, where $\mathbf{p}$ and $\mathbf{q}$ are vectors of positive integers whose elements $\mathbf{q}_{i},\,\mathbf{p}_{i}$ denote the desired synchronization order between the phase $\bm{\Phi}_{i}$ and its connected mean field $\bm{\Psi}_{i}$. However, careful engineering of the structure of $\mathbf{A}_{\bm{\Phi}}$ will most likely be needed to minimize competition of different synchronization orders between the links. Finally, $\bm{\gamma}$ is a phase-lag vector (treated as another hyperparameter) whose role is akin to that of the Sakaguchi-Kuramoto model in promoting asymmetric communication[21].
The scalar global order parameter can be defined similarly, $R(t)\in[0,1]$, which is the degree of global synchrony across all the link phases:
@@
Finally, we discuss the vector hyperparameter of slow natural frequencies $\bm{\omega}_{0}$. While its distribution and values are not constrained a priori, we consider the simple binary case where a set percentage of links share the same natural frequency $\omega_{0}\neq 0$ (a hyperparameter), and the rest are set to 0. We denote the density of links with nonzero natural frequencies as $\Lambda(\omega_{0})\in(0,1]$, another hyperparameter.
@@
We note that we assume a 1-to-1 correspondence between the phase of link strength modulation and the phases of the links themselves. Therefore, the synchronization of the glial integrators is directly reflected in the neuronal dynamics as synchronized link strength modulations. While this was done mainly for simplicity, it was also experimentally observed that the oscillations of the type of chemical signals present in astrocytes may be 1:1 synchronized to the oscillations of the sort of biomechanical oscillations that are found in astrocyte hotspots[9]. However, $f$ can take on other forms representing how the chemical dynamics of the astrocytes are translated to these mechanical waves.

### v4

# Rhythmic sharing: A bio-inspired paradigm for zero-shot adaptive learning in neural networks

@@ Abstract
The brain rapidly adapts to new contexts and learns from limited data, a coveted characteristic that artificial intelligence (AI) algorithms struggle to mimic. Inspired by the mechanical oscillatory rhythms of neural cells, we developed a learning paradigm utilizing link strength oscillations, where learning is associated with the coordination of these oscillations. Link oscillations can rapidly change coordination, allowing the network to sense and adapt to subtle contextual changes without supervision. The network becomes a generalist AI architecture, capable of predicting dynamics of multiple contexts including unseen ones. These results make our paradigm a powerful starting point for novel models of cognition. Because our paradigm is agnostic to specifics of the neural network, our study opens doors for introducing rapid adaptive learning into leading AI models.
@@ 1.1 A new learning paradigm
The second part of the paradigm is that the phases of each link can change depending on the information it processes, e.g., different states, as illustrated in fig. 1C. We propose that links that are active in processing a given state can adjust their phases rapidly to synchronize to other links. This synchronization is mediated by objects that can integrate information from subsets of links (inspired by the integrative role of astrocytes and denoted as stars in fig. 1C).
Because input of a different state results in active link changes, state changes manifest as collective changes to synchronization of the link phases as illustrated in fig. 1C. Therefore, link phase synchronization also functions as a classification token of each state.
@@ 2.1.3 Detecting warning signs of catastrophic phenomena
A well-known chaotic system that exhibits this behavior is the 3D Lorenz system[15]. Its internal parameter $\lambda_{\text{Lorenz}}$ is responsible for state changes. The system exhibits a chaotic trajectory when $\lambda_{\text{Lorenz}}=24.5$. When $\lambda_{\text{Lorenz}}$ abruptly changes to $23.5$, the chaotic trajectory shifts into another similar, but transient chaotic trajectory until it suddenly collapses into a dead signal after some probabilistic time[16]. Indeed, fig. 3C shows the close similarity in the spectral properties of the two chaotic trajectories; to resolve the spectral peak differences ($\approx$ 0.1-0.2 Hz), at least 50-100 cycles of the frequency bands containing the peaks must be observed.
@@ 4.2 Rhythmic sharing
where $\circ$ denotes the Hadamard product. $\bm{\Phi}$ is a vector of phases $[\phi_{1}\,...\,\phi_{N_{l}}]^{\text{T}}$ with $N_{l}$ being the total number of links. $\bm{\omega}_{0}$ is a vector of initially imposed natural frequencies of the links. The vector of local mean fields $\bm{\Psi}(t)\in\mathbb{R}^{N_{l}}$ that each link is coupled to is defined by
@@
where $\mathbf{r}\in\mathbb{R}^{N_{l}}$ is a vector of local order parameters whose elements describe the degree of synchronization across the particular local mean field that each link is coupled to. $e^{i\cdot}$ is taken to operate on $\cdot$ element-wise, e.g., $e^{i\bm{\Psi}(t)}=(e^{i\psi_{1}},e^{i\psi_{2}},...,{e^{i\psi_{N_{l}}})}$. $\mathbf{\hat{A}}_{{\bm{\Phi}}_{ij}}\equiv\mathbf{A}_{{\bm{\Phi}}_{ij}}/|\mathbf{A}_{{\bm{\Phi}}_{i\bullet}}|$, where $\mathbf{A}_{\bm{\Phi}}\in\mathbb{R}^{N_{l}\times N_{l}}$ is a random phase adjacency matrix, binary for simplicity, whose elements equal 1 if the links are connected and 0 otherwise, and the normalization factor $|\mathbf{A}_{{\bm{\Phi}}_{i\bullet}}|$ is the $L^{1}$ norm of the $i$-th row of $\mathbf{A}_{\bm{\Phi}}$. The density of non-zero elements of $\mathbf{A}_{\bm{\Phi}}$ matrix is a tuneable hyperparameter. We note in passing that while we limit the focus of this study to 1:1 synchronization, arbitrary orders of $p:q$ synchronization (for clusters) may be readily implemented by modifying the argument of the right-hand side of the eq. 3 by taking $\sin(\bm{\Psi}-\bm{\Phi})\rightarrow\sin(\mathbf{p}\circ\bm{\Psi}-\mathbf{q}\circ\bm{\Phi})$, where $\mathbf{p}$ and $\mathbf{q}$ are vectors of positive integers whose elements $\mathbf{q}_{i},\,\mathbf{p}_{i}$ denote the desired synchronization order between the phase $\bm{\Phi}_{i}$ and its connected mean field $\bm{\Psi}_{i}$. However, careful engineering of the structure of $\mathbf{A}_{\bm{\Phi}}$ will most likely be needed to minimize competition of different synchronization orders between the links. Finally, $\bm{\gamma}$ is a phase-lag vector (treated as another hyperparameter) whose role is akin to that of the Sakaguchi-Kuramoto model in promoting asymmetric communication[21].
Finally, we discuss the vector hyperparameter of slow natural frequencies $\bm{\omega}_{0}$. While its distribution and values are not constrained a priori, we consider the simple binary case where a set percentage of links share the same natural frequency $\omega_{0}\neq 0$, itself a hyperparameter, and the rest are set to 0. We denote the density of links with nonzero natural frequencies as $\Lambda(\omega_{0})\in(0,1]$, which is also a hyperparameter.
The scalar global order parameter used to track the coordination of link phases is defined similarly, $R(t)\in[0,1]$, which is the degree of global synchrony across all the link phases:
@@
We note that we assume a 1-to-1 correspondence between the phase of link strength modulation and the phases of the links themselves. Therefore, the synchronization of the glial mediators is directly reflected in the neuronal dynamics as synchronized link strength modulations. While this was done mainly for simplicity, it was also experimentally observed that the oscillations of the type of chemical signals present in astrocytes may be 1:1 synchronized to the oscillations of the sort of biomechanical oscillations that are found in astrocyte hotspots[9]. However, $f$ can take on other forms representing how the chemical dynamics of the astrocytes are translated to these mechanical waves.